I need help with these - PLEASE!!!!
okay
1. log4 7 = x ( the 4 should be lower - you know!!!)
2. log0.8 10 = x (same thing)
3. logx 343 = 3
Hope these make sense
Thanks
Llo

Answer by stanbon(48569)

1. log4 7 = x
x = [log7] / [log4] = 1.403677461...
---------------------
2. log0.8 10 = x
x = [log10] / [log0.8] = 1/(-0.09691013...)= -10.31885116...
-----------------------
3. logx 343 = 3
x^3 = 343
x = 343^(1/3)
x = 7
===============
Cheers,
Stan H.